Report on the DVI Driver Standard

Joachim Schrod

Secretary

TUG DVI Driver Standards Committee

The DVI Driver Standard will be available in several stages. The basic stage is now called level 0. It covers only those driver capabilities which are really necessary to output a DVI document on an output device. All other driver capabilities will be called features (and may even be realized outside a driver). In the future we will publish several additional standard documents which will cover ranges of features; those documents will represent “tiers” built upon level 0 or on previous tiers. In this way they will be available as future stages of a complete standard. (One may doubt whether the standard will ever be complete as there may be always new features to standardize.)

The basic stage, level 0, consists of three parts:

(1) The pure standard document telling what a driver must be able to do.

(2) Definitions of all file formats spoken of in part 1.

(3) A rationale describing why the committee has chosen the given definition in part 1, recalling discussions that led to particular decisions.

A draft of the level 0 document is about to be published for public review. Part 1 of the draft is (almost) ready; a few spelling errors and such have to be removed. Part 2 was ready, but D. E. Knuth has changed the GF documentation, and this change must be incorporated. Part 3 exists only in part.

The committee will publish the draft as soon as possible. It may be that the draft of the rationale will not be finished in time; in that event we will publish part 1 by itself. This is considered to be useful (although not desirable) so that we will get responses very soon — and especially to change the status from “draft” to “released” as soon as possible. The file formats will not be published in TUGboat; they are available on several file servers. For people who do not have access to file servers I’ve prepared a brochure covering all file formats.

When complete, the standard will be published in the TEXniques series. The style will be modified slightly to follow formal standards conventions. The body of the standard will form the main text; this will be followed by a number of “annexes”. The file formats will come immediately after the main text, as “normative” annexes; that is, these format specifications are an integral part of the standard, but the presentation of each is self-contained and

TUGboat, Volume 0 (2001), No. 0  preliminary draft, 15 May 1995 17:21  1001

too large to be appropriate in the main text. Finally, the rationale will appear as an “informative” annex, to present information that is not an integral part of the standard, but is intended to help a user in understanding it.

Future work

What tier will come next, i.e., what driver feature will be looked at next, is still unclear. There is public pressure to tackle the area of graphics inclusion at an early date; others want to touch areas such as page selection, etc., first. So this remains an open problem. We invite all parties to bring proposals to the committee. My personal opinion is that a proposal for a new tier received early will be handled early. So if someone is eager to see a specific topic addressed, he or she should do work on this topic and send us the result of the work. (We will be glad to acknowledge contributors.)

⋄ Joachim Schrod Technical University of Darmstadt Institut für Theoretische

Informatik Alexanderstraße 10 W-6100 Darmstadt Federal Republic of Germany

                                                                     Bitnet: xitijsch@ddathd21

preliminary draft, 15 May 1995 17:21 preliminary draft, 15 May 1995 17:21